# UniRGB-IR: A Unified Framework for RGB-Infrared Semantic Tasks via Adapter Tuning

### v2

@@ IV-A RGB-IR Object Detection
TABLE V: Ablation studies of key components on the FLIR and PST900 datasets. The best results are highlighted in bold.
@@
IV-B RGB-IR Semantic Segmentation
Datasets. Our semantic segmentation experiments are performed on the two public RGB-IR semantic segmentation datasets: PST900 [62] and MFNet [18]. The PST900 dataset is divided into 597 pairs for training and 288 pairs for testing, containing five categories (background, fire extinguisher, backpack, hand drill, and survivor). The MFNet dataset provides 1,569 pairs and contains nine classes (Car, Curve, Person, Guardrail, Car stop, Bike, Bump, Color cone, and Background). The dataset is divided into three parts: training, verification and testing according to the ratio of 2:1:1.
Settings. Similarly, as the RGB-IR object detection task, we incorporate our method into the SETR [81] basic framework and implement it on the MMSegmentation library. The fine-tuning process spins a total of 10K iterations with an initial learning rate of 0.01. We employ the SGD optimizer and set the batch size to 16.
Results. The quantitative results of the different RGB-IR segmentation methods on the PST900 and MFNet dataset are shown in Table III and Table IV respectively. The comparison results show that our model significantly outperforms other methods. Specifically, our model obtains the best performance in terms of both the mACC and mIoU. In PST900 dataset, our model performs competitive performance in the Backpack, Hand-Drill and Survivor, outperforming the second-best methods by 2.3%, 0.6% and 0.4% IoU, respectively. In predicting Color Cone class on MFNet dataset, our model achieves the first and second best in terms of both evaluation metrics, with about 11% Acc higher than that of the second-best. These results show that our UniRGB-IR can also be used flexibly for segmentation tasks without any modification, which validates the universality of our method.
TABLE VI: Ablation of adding SFI module to different stages.
@@ IV-B RGB-IR Semantic Segmentation
TABLE VII: Ablation of the different attention mechanisms.
@@
Global Attention [82] Quadratic 36.1 72.8 30.6
Pale Attention [42] Linear 31.6 68.4 24.6
Deformable Attention [43] Linear 44.1 81.4 40.2
IV-C Ablation Study
Ablation for components. To investigate the contribution of the SFI and MFP modules, we gradually add each module to the baseline model as shown in Table V. We stack RGB and IR images and feed them into the Cascade R-CNN detector and SETR framework (ViT backbone) for full fine-tuning on the corresponding tasks, which are considered as our baseline models. Then, we freeze the ViT model and introduce the MFP module to extract contextual multi-scale features (element-wise addition), which results in 1.6% mAP and 1.3% mIoU improvements. Finally, we replace the element-wise addition operation with the SFI module and further improve the mAP and mIoU metrics by 2.9% and 4.0% respectively, which achieve the best performance on both datasets.
SFI module at different stages. We add the SFI module to the ViT pre-trained model at the beginning of different stages. From Table VI, we can find that by adding the SFI module at the first stage, the detector achieves 41.7% mAP on FLIR dataset. After adding the SFI module in the second and third stages, the performance further improved by about 2% and 3% mAP, respectively. However, continuing to add it to the final stage will reduce the detection performance while also increasing computational overhead. Therefore, we add the SFI module from the first stage to the third stage of ViT model.
Attention type in SFI module. Since the attention mechanism in our SFI module is replaceable, we adopt three popular attention mechanisms in our UniRGB-IR to discuss their impact on model performance. As shown in Table VII, the detector achieves the best performance with linear complexity by utilizing deformable attention. Thus, the deformable attention is more suitable for our framework and utilized as the default configuration. It is worth noting that it can be replaced by other attention mechanisms to further achieve superior performance.
Figure 6: Visualization of intermediate results. The $\boldsymbol{F_{mfp}}$ and $\boldsymbol{F_{sfi}}$ features from the first stage are visualized in the third and fourth columns. The tSNE visualizations are also shown in the last two columns.
Figure 7: Training efficiency analysis on the FLIR dataset.
IV-D Visualization Analysis
[…]
TABLE VIII: Quantitative comparison of RGB-IR salient object detection task on VT821, VT1000 and VT5000 datasets. * represents RGB-D SOD transformed into RGB-T SOD. The best results are highlighted in red and the second-best are highlighted in blue.
Model VT821 VT1000 VT5000
[…]
MMCI* [83] 0.763 0.784 0.618 0.087 0.886 0.892 0.803 0.039 0.827 0.859 0.714 0.055
TANet* [84] 0.818 0.852 0.717 0.052 0.902 0.912 0.838 0.030 0.847 0.883 0.754 0.047
S2MA* [85] 0.811 0.813 0.709 0.098 0.918 0.912 0.848 0.029 0.853 0.864 0.743 0.053
JLDCF* [86] 0.839 0.830 0.726 0.076 0.912 0.899 0.829 0.030 0.861 0.860 0.739 0.050
MTMR [87] 0.725 0.815 0.662 0.109 0.706 0.836 0.715 0.119 0.680 0.795 0.595 0.114
M3S-NIR [88] 0.723 0.859 0.734 0.140 0.726 0.827 0.717 0.145 0.652 0.780 0.575 0.168
SGDL [89] 0.765 0.847 0.731 0.085 0.787 0.856 0.764 0.090 0.750 0.824 0.672 0.089
FMSF [90] 0.760 0.796 0.640 0.080 0.873 0.899 0.823 0.037 0.814 0.864 0.734 0.055
MIDD [91] 0.871 0.895 0.803 0.033 0.915 0.933 0.880 0.027 0.868 0.896 0.799 0.043
ADF [92] 0.810 0.842 0.717 0.077 0.910 0.921 0.847 0.034 0.864 0.891 0.778 0.048
LSNet [93] 0.877 0.911 0.827 0.033 0.924 0.936 0.887 0.022 0.876 0.916 0.827 0.036
Figure 8: Saliency maps of proposed UniRGB-IR and seven representative SOTA RGB-IR SOD methods in difference scenes. From top to bottom are visualization examples from the VT821, VT1000 and VT5000 datasets respectively.
V Task Extension
V-A RGB-IR Salient Object Detection
Datasets. RGB-IR salient object detection (SOD) experiments are performed on the three public datasets: VT821 [87], VT1000 [89] and VT5000 [92]. The VT821 dataset includes 821 registered RGB and IR images. The VT1000 dataset contains 1000 registered RGB-IR images with simple scenes and aligned images. The VT5000 dataset is a recent large-scale RGB-IR dataset, including a full-day scene under various limited light conditions. As usual in [92], we utilize 2500 image pairs in the VT5000 dataset as the training dataset, and the remaining image pairs along with the image pairs from the VT821 and VT1000 datasets are used as the test datasets.
@@ V-A RGB-IR Salient Object Detection
Quantitative Comparison. Table VIII reports the quantitative comparison results. As can be seen from Table VIII, our UniRGb-IR outperforms SOTA methods both on VT1000 and VT5000 datasets in all evaluation metrics. Specifically, the $S$, $adpE$, $adpF$ and $MAE$ matrics of our UniRGB-IR achieve 0.906, 0.935, 0.849 and 0.027 on VT5000, all of which are higher than the previous competitor LSNet [93]. These remarkable results clearly indicate that the saliency maps predicted by UniRGB-IR are close to the corresponding ground-truths, which verifies the effectiveness of our method.
Qualitative Evaluation. Qualitative RGB-IR SOD comparisons are shown in Figure 8. The first and third rows show scenes containing small objects. The second and fifth rows show the salient object with the complex backgrounds. All these scenes are suitably handled by our UniRGB-IR, which verifies that UniRGB-IR is more effective while providing high-quality results with appropriate detection capabilities.
V-B RGB-IR Face Recognition
Datasets. The face recognition task is performed on the IRIS thermal/visible face dataset [94], which contains pairs of thermal and visible face images captured simultaneously under different illumination conditions and pose rotations as shown in Figure 9. The dataset has a total of 4228 pairs of thermal and visible images from 30 subjects. We randomly select 80% of the images as the training set and the rest as the test set to verify our method.
Metrics. For RGB-IR face recognition task, we utilize two metrics to evaluate the recognition results, called Rank-1 accuracy and True Positive Rate (TPR) at False Acceptance Rate (FAR) of 1% and 0.1%. Rank-1 accuracy is defined as the proportion of correctly predicted face image pairs (True positive) among the total number of image pairs and TPR at an FAR is described as the number of correct predictions at a specific percentage of acceptable false predictions.
Settings. We conduct the RGB-IR face recognition experiments based on the Insightface [95] library and incorporate our method into the framework to replace the backbone network. During training, the images are resized to a size of $224\times 224$ pixels and the labels contain class information only. At this stage, the categorical cross-entropy loss is used for multimodal face recognition. The training process is terminated at 20 epochs and the learning rate is initialized as 1$\times$10-4 with SGD optimizer and divided by 10 at 10, 14, 17 epochs.
Figure 9: Example of visible and infrared images from IRIS face dataset. Each subject in the dataset contains different pose rotations and illumination conditions.
TABLE IX: Quantitative comparison of RGB-IR face recognition task on IRIS thermal/visible face dataset.
Methods Modality Rank-1(%) TPR@FAR = 1% TPR@FAR = 0.1%
Arcface [95] IR 70.4 80.2 53.9
Adaface [96] IR 76.3 85.1 62.3
Arcface [95] RGB 61.6 75.7 49.2
Adaface [96] RGB 64.6 79.6 52.7
Arcface+UniRGB-IR (Ours) RGB+IR 83.3 88.5 65.8
Adaface+UniRGB-IR (Ours) RGB+IR 84.2 89.4 66.2
Results. Table IX presents the rank-1 identification and verification rates on IRIS thermal/visible face dataset. It can be seen from the table that the recognition results in the IR modality are obviously better than those in the RGB modality. Besides, after using our UniRGB-IR framework, the recognition performance is further improved. Specifically, Our UniRGB-IR with Adaface achieves 84.3%, 89.4% and 66.2% in Rank-1, TPR@FAR = 1% and TPR@FAR = 0.1% respectively, which are the highest among these results.
VI Conclusion
@@
In this paper, we proposed an efficient and universal framework (named UniRGB-IR) for RGB-IR semantic tasks. Our framework contains a Multi-modal Feature Pool module and a Supplementary Feature Injector module. The former extracts contextual multi-scale features from two modality images, and the latter adaptively injects the features into the transformer model. These two modules can be efficiently optimized to complement the pre-trained foundation model with richer RGB-IR features. To evaluate the effectiveness of our method, we incorporated the ViT-Base model into the framework as the pre-trained foundation model and performed various RGB-IR downstream tasks. Extensive experiments verify that our UniRGB-IR can be effectively leveraged as a unified framework for RGB-IR semantic tasks. We believe that our framework holds potential for expansion to a broader range of multi-modal tasks.

### v4

@@ 4.2. RGB-IR Semantic Segmentation
Table 5. Ablation studies of key components on the FLIR and PST900 datasets. The best results are highlighted in bold.
@@
Settings. Similarly, as the RGB-IR object detection task, we incorporate our method into the SETR (Zheng et al., 2021) basic framework and implement it on the MMSegmentation library. The fine-tuning process spins a total of 10K iterations with an initial learning rate of 0.01. We employ the SGD optimizer and set the batch size to 16.
Table 6. Ablation of adding SFI module to different stages.
@@
Table 7. Ablation of the different attention mechanisms.
@@
Global Attention (Vaswani et al., 2017) Quadratic 36.1 72.8 30.6
Pale Attention (Wu et al., 2022c) Linear 31.6 68.4 24.6
Deformable Attention (Zhu et al., 2020) Linear 44.1 81.4 40.2
Results. The quantitative results of the different RGB-IR segmentation methods on the PST900 dataset are shown in Table 3. The comparison results show that our model significantly outperforms other methods. Specifically, our model obtains the best performance in terms of both the mACC and mIoU. Besides, our model performs competitive performance in the Backpack, Hand-Drill and Survivor, outperforming the second-best methods by 2.3%, 0.6% and 1.1% IoU, which strongly demonstrates the effectiveness of our UniRGB-IR.
4.3. RGB-IR Salient Object Detection
Datasets. Our salient object detection (SOD) experiments are performed on the three public datasets: VT821 (Wang et al., 2018), VT1000 (Tu et al., 2019b) and VT5000 (Tu et al., 2022). The VT821 dataset includes 821 registered RGB and IR images. The VT1000 dataset contains 1000 registered RGB-IR images with simple scenes and aligned images. The VT5000 dataset is a recent large-scale RGB-IR dataset, including a full-day scene under various limited light conditions. As usual in (Tu et al., 2022), we utilize 2500 image pairs in the VT5000 dataset as the training dataset, and the remaining image pairs along with the image pairs from the VT821 and VT1000 datasets are used as the test datasets.
@@ 4.3. RGB-IR Salient Object Detection
Results. Table 4 reports the quantitative comparison results. As can be seen from Table 4, our UniRGB-IR outperforms SOTA methods both on VT1000 and VT5000 datasets in all evaluation metrics. Specifically, the $S$, adpEadpE, adpFadpF and $MAE$ matrics of our UniRGB-IR achieve 0.906, 0.935, 0.849 and 0.027 on VT5000, all of which are higher than the previous competitor UniT (Guo et al., 2024). These remarkable results clearly indicate that the saliency maps predicted by UniRGB-IR are close to the corresponding ground-truths.
Figure 5. Visualization of intermediate results. The $\boldsymbol{F_{mfp}}$ and $\boldsymbol{F_{sfi}}$ features from the first stage are visualized in the third and fourth columns. The tSNE visualizations are also shown in the last two columns.
4.4. Ablation Study
Ablation for components. To investigate the contribution of the SFI and MFP modules, we gradually add each module to the baseline model as shown in Table 5. We stack RGB and IR images and feed them into a frozen standard ViT model equipped with the Cascade R-CNN detection head and SETR segmentation head. The results obtained by fine-tuning the patch embedding layer and the detection or segmentation head are considered as our baseline. Then, we also freeze the ViT model and introduce the MFP module to extract contextual multi-scale features (element-wise addition), which results in 2.7% mAP and 3.0% mIoU improvements. Finally, we replace the element-wise addition operation with the SFI module and further improve the mAP and mIoU metrics by 2.9% and 4.0% respectively, which achieve the best performance on both datasets.
SFI module at different stages. Since we utilize the entire standard ViT encoder as the foundation model, we perform ablation experiments by adding the SFI module to the ViT pre-trained model at the beginning of different stages. From Table 6, we can find that by adding the SFI module at the first stage, the detector achieves 41.7% mAP on FLIR dataset. After adding the SFI module in the second and third stages, the performance further improved by about 2% and 3% mAP, respectively. However, continuing to add it to the final stage will reduce the detection performance while also increasing computational overhead. Therefore, we add the SFI module from the first stage to the third stage of ViT model.
Attention type in SFI module. Since the attention mechanism in our SFI module is replaceable, we adopt three popular attention mechanisms in our UniRGB-IR to discuss their impact on model performance. As shown in Table 7, the detector achieves the best performance with linear complexity by utilizing deformable attention. Thus, the deformable attention is more suitable for our framework and utilized as the default configuration. It is worth noting that it can be replaced by other attention mechanisms to further achieve superior performance.
Figure 6. Training efficiency analysis on the FLIR dataset.
4.5. Visualization Analysis
Intermediate results. To illustrate the effectiveness of the SFI module, we visualize the intermediate results on the FLIR dataset. From $F_{mfp}$ and $F_{sfi}$ in Figure 5, we can find that the foreground objects in the $F_{sfi}$ become salient through the SFI module. Furthermore, we also visualize the tSNE maps of $F_{mfp}$$F_{vit}$ and $F_{sfi}$$F_{vit}$ respectively. After using the SFI module, the distribution of injected features $F_{sfi}$ is more concentrated than the distribution of ViT features $F_{vit}$, indicating that the required richer RGB-IR features can be well supplemented into the ViT model through the SFI module.
5. Conclusion
@@
In this paper, we proposed an efficient and scalable framework (named UniRGB-IR) for RGB-IR semantic tasks. The framework contains a Multi-modal Feature Pool module and a Supplementary Feature Injector module. The former extracts contextual multi-scale features from two modality images, and the latter adaptively injects the features into the transformer model. These two modules can be efficiently optimized to complement the pre-trained foundation model with richer RGB-IR features. To evaluate the effectiveness of our method, we incorporated the ViT-Base model into the framework as the pre-trained foundation model and performed various RGB-IR semantic tasks. Extensive experiments verify that our UniRGB-IR can be effectively leveraged as a unified framework for RGB-IR downstream tasks. We believe that our method can be applied to more multi-modal real-world applications.
Acknowledgements.
This work was supported by the Fundamental Research Funds for the Central Universities.